Two spherical objects have masses of 7.5 x 108 kg and 6.5 x 108 kg. Their centers are separated by a distance of 2.5 x 106 m.


1
Expert's answer
2022-04-07T13:10:06-0400

The gravitation force between spheres

"F=G\frac{m_1m_2}{r^2}"

"F=6.67*10^{-11}*\frac{7.5*10^8*6.5*10^8}{(2.5*10^6)^2}=5.2*10^{-6}\:\rm N"